# MiniCheck: Efficient Fact-Checking of LLMs on Grounding Documents

### v1

@@ Abstract
Recognizing if LLM output can be grounded in evidence is central to many tasks in NLP: retrieval-augmented generation, summarization, document-grounded dialogue, and more. Current approaches to this kind of “fact-checking” are based on verifying each piece of a model generation against potential evidence using an LLM. However, this process can be very computationally expensive, requiring many calls to LLMs to check a single response. In this work, we show how to build small models that have GPT-4-level performance but for 400x lower cost. We do this by constructing synthetic training data with GPT-4, which involves creating realistic yet challenging instances of factual errors via a structured generation procedure. Training on this data teaches models to check each fact in the claim and recognize synthesis of information across sentences. For evaluation, we unify pre-existing datasets into a benchmark LLM-AggreFact, collected from recent work on fact-checking and grounding LLM generations. Our best system MiniCheck-FT5 (770M parameters) outperforms all systems of comparable size and reaches GPT-4 accuracy. We release LLM-AggreFact, code for data synthesis, and models. [^1] Available at https://github.com/Liyan06/MiniCheck.
@@ 1 Introduction
Past work has largely dealt with these problems separately. When post-hoc verifying the content of closed-book generated answers, a separate retrieval stage is needed Gao et al. (2023); Malaviya et al. (2024); Jacovi et al. (2024), and many statements cannot be sourced or may have conflicting information available Wang et al. (2023); Glockner et al. (2024). Grounding statements in these retrieved documents adds another layer of difficulty. In contrast, when the grounding is known in settings like summarization, the attribution problem is often framed as document-level textual entailment Nie et al. (2020); Yin et al. (2021)  and has been studied extensively for smaller language models Falke et al. (2019); Goyal and Durrett (2020, 2021); Laban et al. (2022); Tang et al. (2023a).
The problems in this space have a shared primitive operation: the need to check a statement against grounding documents, either retrieval-augmented content or post-hoc retrieved evidence. We call this primitive fact-checking on grounding documents, shown in Figure 1. Implementations of this primitive need to be accurate, spotting subtle errors while maintaining a low false positive rate, as most statements are usually correct. They also need to be efficient: a single LLM response may contain dozens of facts to verify, and self-verification with an LLM may increase cost by an order of magnitude Weng et al. (2023); Gero et al. (2023). For instance, the 110-150 word biographies in FActScore (Min et al., 2023) contain 26-41 atomic facts that are checked against 5 documents each, resulting in 130-205 entailment checks.
In this work, we build an efficient system for fact-checking on grounding documents. Our key insight is to develop a new synthetic training dataset which is tailored to the complexities of the fact-checking task. Unlike standard distillation from LLMs Taori et al. (2023); Hsieh et al. (2023), this setting differs in that we do not necessary have access to task instances that we can label with strong LLMs. For instance, in datasets like ExpertQA Malaviya et al. (2024), even the inputs to the LLM are expert-written questions. As a result, we synthesize challenging fact-checking instances from the ground up, as a scalable way to teach a small model how to simultaneously verify multiple facts in a sentence against multiple sentences in grounding documents. Our system, MiniCheck, is an instance of Flan-T5 Chung et al. (2022) fine-tuned on this data plus standard entailment data Nie et al. (2020).
@@ 3 Methodology: Training Data Synthesis
To address these challenges, new data is required. Existing datasets like MNLI Williams et al. (2018) and ANLI Nie et al. (2020) do not feature instances that reflect the complexity of LLM fact-checking. Annotation of real errors is challenging to scale; datasets of such errors (including those in LLM-AggreFact) are largely test-only.
Our goal is to construct a dataset $\{(D_{i},c_{i},y_{i})\}_{i=1}^{N}$ of $N$ instances of documents $D_{i}$ paired with claims $c_{i}$ with label $y_{i}\in\{0,1\}$, using two novel synthetic data generation methods (Figure 3). Statistics about our final synthetic training data can be found in Table 1. Additional details, including the sources of claims and documents and examples of generated data, can be found in Appendix D. We provide all prompts and quality assurance details in Appendix H.
@@ 3.3 MiniCheck Models
As models trained on the ANLI dataset Nie et al. (2020) have demonstrated strong performance Kamoi et al. (2023); Honovich et al. (2022), we integrate our data with the ANLI dataset for fine-tuning deberta-v3-large He et al. (2021) and flan-t5-large Chung et al. (2022). We take a subset (21K) of the ANLI training data, selecting examples where their trained entailment models made incorrect predictions during dataset construction. Training on more of ANLI was not effective.
@@ 4.1 Benchmark Details
Table 2: Performance (BAcc) of models on the test set of LLM-AggreFact without per-dataset threshold tuning. Models are split into LLM-based fact-checkers | specialized fact-checkers | Ours. We highlight the best performance for each dataset, where multiple green highlights indicate systems indistinguishable from the best according to a paired bootstrap test with p-value < 0.05. Details for -Dectx and -Decmp are in Section 7. Our MiniCheck models outperform other specialized evaluators and MiniCheck-FT5 reaches the performance of GPT-4.
@@ 9 Conclusion
The effectiveness of our synthetic data is demonstrated by the improved performance when training on it across various model architectures. As LLMs continue to advance, the quality of the synthetic data generated using our approach is also expected to improve. However, we did not conduct a formal human evaluation to evaluate the quality of our synthetically generated data (i.e., how many instances are mislabeled).
@@ A.1 Intrinsic Evaluation on C2D/D2C
Figure 5 shows the in-distribution performance of FT5-C2D and FT5-D2C as optimal performance in each sub-figure, represented by the black dashed lines. We observe that models trained on synthetic data with simplified construction steps (C2D-Simp and D2C-Simp) fail to develop the desired properties we expect. FT5-D2C-S performs even worse than random chance on the held-out set of C2D. In contrast, models trained on C2D and D2C outperform all other fact-checkers on the OOD held-out set of D2C and C2D, respectively. Additionally, we note that the model trained on 163K ANLI data points fail to reach the performance of models trained solely on 7K synthetic data. These findings demonstrate that both of our synthetic data generation methods can effectively encourage models to pay more attention to multiple atomic facts and reason over multiple sentences, even with a limited amount of training data.
A.2 Ablation of D2C/C2D
Appendix B Additional Results
B.1 Results using Threshold Tuning
@@ Appendix D Synthetic Data Details
In our C2D method, we choose around 400 claims from Wikipedia that have cited web articles Kamoi et al. (2023); Petroni et al. (2023) to generate synthetic documents. In our D2C method, we scraped around 300 Google News articles since November 2023 from diverse topic categories, including science, politics, world, entertainment, business, and technology. Each document is approximately 500 words. Statistics of our generated data can be found in Table 1. See Appendix H.1 for how we maintain the quality of our synthetic data.
@@
Examples of synthetic data for C2D and D2C can be found in Table 14 and 15 .
@@ Appendix G Training Details
Zero-shot factual consistency evaluation Tabel 22
@@ H.1 Quality Assurance for Generations
- Lord Steven Regal wan the World Television Championship.
- The Nasty Boys wan and the World Tag Team Championship.
@@
Your task is to generate a pair of sentences so that the provided claim can be entailed by the sentence pair. You must make sure that the claim can only be deducted by combining the information from the two sentences that contain unique information.
@@
We are creating a news article (one paragraph) in the style of The New York Time. We will give you a list of facts to use when writing your article. You must include all the facts in the list. Never state deduced facts or conclusions. The article should stick to the fact list pretty closely. Include as many sentences as needed to write each fact from the list of facts.

### v2

@@ Abstract
Recognizing if LLM output can be grounded in evidence is central to many tasks in NLP: retrieval-augmented generation, summarization, document-grounded dialogue, and more. Current approaches to this kind of fact-checking are based on verifying each piece of a model generation against potential evidence using an LLM. However, this process can be very computationally expensive, requiring many calls to a model to check a single response. In this work, we show how to build small fact-checking models that have GPT-4-level performance but for 400x lower cost. We do this by constructing synthetic training data with GPT-4, which involves creating realistic yet challenging instances of factual errors via a structured generation procedure. Training on this data teaches models to check each fact in the claim and recognize synthesis of information across sentences. For evaluation, we unify datasets from recent work on fact-checking and grounding LLM generations into a new benchmark, LLM-AggreFact. Our best system MiniCheck-FT5 (770M parameters) outperforms all systems of comparable size and reaches GPT-4 accuracy. We release LLM-AggreFact, code for data synthesis, and models.111 https://github.com/Liyan06/MiniCheck; see latest leaderboard at llm-aggrefact.github.io.
@@ 1 Introduction
Past work has largely dealt with these problems separately. We can post-hoc verify closed-book generated answers by retrieving supporting documents and checking the answer against them Gao et al. (2023); Malaviya et al. (2024); Jacovi et al. (2024), which requires precise checking as many statements are not exactly supported or may have conflicting information available Wang et al. (2023); Glockner et al. (2024). When the grounding is known in settings like summarization, the attribution problem can be cleanly framed as document-level textual entailment Nie et al. (2020a); Yin et al. (2021)  and has been studied extensively for smaller language models Falke et al. (2019); Goyal and Durrett (2020, 2021); Laban et al. (2022); Tang et al. (2023a).
The problems in this space have a shared primitive operation: the need to check a statement against grounding documents, either retrieval-augmented content or post-hoc retrieved evidence. We call this primitive fact-checking on grounding documents, shown in Figure 1. Implementations of this primitive need to be accurate, spotting subtle errors while maintaining a low false positive rate, as most generated statements are correct. They also need to be efficient: a single LLM response may contain dozens of facts to verify, and self-verification with an LLM may increase cost by an order of magnitude Weng et al. (2023); Gero et al. (2023). For instance, the 110-150 word biographies in FActScore (Min et al., 2023) contain 26-41 atomic facts that are checked against 5 documents each, resulting in 130-205 entailment checks.
In this work, we build an efficient system for fact-checking on grounding documents. Our key insight is to develop a new synthetic training dataset which is tailored to the complexities of the fact-checking task. Unlike standard distillation from LLMs Taori et al. (2023); Hsieh et al. (2023), this setting differs in that we do not necessary have access to task instances that we can label with strong LLMs. For instance, in datasets like ExpertQA Malaviya et al. (2024), even the inputs to the LLM are expert-written questions. As a result, we synthesize challenging fact-checking instances from the ground up, as a scalable way to teach a small model how to simultaneously verify multiple facts in a sentence against multiple sentences in grounding documents. Our system, MiniCheck, is an instance of Flan-T5 Chung et al. (2022) fine-tuned on this data plus standard entailment data Nie et al. (2020a).
@@ 3 Methodology: Training Data Synthesis
To address these challenges, new data is required. Existing datasets like MNLI Williams et al. (2018) and ANLI Nie et al. (2020a) do not feature instances that reflect the complexity of LLM fact-checking. Annotation of real errors is challenging to scale; datasets of such errors (including those in LLM-AggreFact) are largely test-only.
Our goal is to construct a dataset $\{(D_{i},c_{i},y_{i})\}_{i=1}^{N}$ of $N$ instances of documents $D_{i}$ paired with claims $c_{i}$ with label $y_{i}\in\{0,1\}$, using two novel synthetic data generation methods (Figure 3). Statistics about our final synthetic training data can be found in Table 1. A small-scale human evaluation of the synthetic data quality can be found in Appendix A.3. Additional details, including the sources of claims and documents and examples of generated data, can be found in Appendix D. We provide all prompts and quality assurance details in Appendix H.
@@ 3.3 MiniCheck Models
As models trained on the ANLI dataset Nie et al. (2020a) have demonstrated strong performance Kamoi et al. (2023); Honovich et al. (2022), we integrate our data with the ANLI dataset for fine-tuning deberta-v3-large He et al. (2021) and flan-t5-large Chung et al. (2022). We take a subset (21K) of the ANLI training data, selecting examples where their trained entailment models made incorrect predictions during dataset construction. Training on more of ANLI was not effective.
@@ 4.1 Benchmark Details
Table 2: Performance (BAcc) of models on the test set of LLM-AggreFact without per-dataset threshold tuning. Models are split into LLM-based fact-checkers | specialized fact-checkers | Ours. We highlight the best performance for each dataset, where multiple green highlights indicate systems indistinguishable from the best according to a paired bootstrap test with 1000 runs and p-value < 0.05. Details for -Dectx and -Decmp are in Section 7. Our MiniCheck models outperform other specialized evaluators and MiniCheck-FT5 reaches the performance of GPT-4.
@@ 9 Conclusion
The effectiveness of our synthetic data is demonstrated by the improved performance when training on it across various model architectures. We find that using a pair of sentences to support a fact is a simple method that yields useful training data for our models. However, there are other possible strategies for how atomic facts or claims could be expanded into multiple sentences. We believe that constructing more complex and higher-quality data could be a future direction not only for this work but also for other related tasks. As LLMs continue to advance, the quality of the synthetic data generated using our approach is also expected to improve.
@@ A.1 Intrinsic Evaluation on C2D/D2C
Figure 5 shows the in-distribution performance of FT5-C2D and FT5-D2C as optimal performance in each sub-figure, represented by the black dashed lines. We observe that models trained on synthetic data with simplified construction steps (C2D-Simp and D2C-Simp) fail to develop the desired properties we expect. FT5-D2C-S performs even worse than random chance on the held-out set of C2D. In contrast, models trained on C2D and D2C outperform all other fact-checkers on the OOD held-out set of D2C and C2D, respectively. Additionally, we note that the model trained on 163K ANLI data points fail to reach the performance of models trained solely on 7K synthetic data. Our synthetic data generation methods can effectively encourage models to pay more attention to multiple atomic facts and reason over multiple sentences, even with a limited amount of training data.
@@
A.2 Ablation of D2C/C2D
A.3 Human Evaluation of Synthetic Data
While we use automatic entailment checks to ensure label quality in the training data construction pipeline, we conduct a small scale human evaluation to measure the quality of the generated data.
We randomly chose 40 (document, claim) pairs from each of the C2D and D2C training data. Three authors of the work independently annotated these 40 $\times$ 2 datapoints as supported or not supported (without seeing the gold label), with an average annotation time of around 1 min and 40 seconds per example. Among the three annotators, we computed an annotation agreement using Fleiss’ Kappa, obtaining a score of 0.51 for the C2D samples and 0.70 on the D2C samples, indicating moderate to substantial agreement. The annotators adjudicated cases with disagreement to reach a consensus on the final factuality of the labels. We refer to these as the ground truth labels.
Compared to these ground truth labels, the labels given to our training samples have an accuracy of 80% on C2D and 78% on D2C, respectively. We calculated similar accuracy values for the human annotators, and the average across the three annotators’ accuracies was 85% on C2D and 88% on D2C. These accuracies demonstrate that the automatic labels are a bit lower-quality than single-annotator labels on D2C, but close on C2D.
Many of the disagreement cases are classic cases of subjectivity in NLI Pavlick and Kwiatkowski (2019); Chen et al. (2020); Nie et al. (2020b). For example, for the C2D claim “Located in the Scottish Highlands, she lived with and later married James Ballard in Spean Bridge.”, the generated passage only references the couple living in the town of Spean Bridge after the marriage, making it unclear whether or not they got married there. However, this is a reasonable supposition to make. Our results show that our data is useful for training in spite of these subjective examples.
Appendix B Additional Results
B.1 Results using Threshold Tuning
@@ Appendix D Synthetic Data Details
In our C2D method, we choose around 400 claims from Wikipedia that have cited web articles Kamoi et al. (2023); Petroni et al. (2023) to generate synthetic documents. In our D2C method, we scraped around 300 Google News articles since November 2023 from diverse topic categories, including science, politics, world, entertainment, business, and technology. Each document is approximately 500 words. Statistics of our generated data can be found in Table 1 and 12. See Appendix H.1 for how we maintain the quality of our synthetic data.
Label Min. 25% 50% 75% Max.
C2D 1 72 147 182 242 359
0 66 136 171 234 359
D2C 1 85 139 162 186 427
0 84 138 161 186 493
Table 12: Length distribution of generated documents. We use the NLTK package for word tokenizatiopn.
@@
Examples of synthetic data for C2D and D2C can be found in Table 15 and 16.
Data Rejection Rate
Since we use GPT-4 for filtering out low-quality examples in our C2D method, we report the rejection rate at different steps that require entailment checks.
During the atomic fact expansion step (step 2), 6% of the final generated sentence pairs, when combined, could not support the original atomic fact. In the supporting document generation step (step 3), 5% of the final documents failed the entailment check. For the non-supporting document generation step (step 4), 53% of the final documents still supported the claim, and these documents were not included in our constructed data. These filtering steps are crucial for improving the training dataset’s quality.
@@ Appendix G Training Details
Zero-shot factual consistency evaluation Tabel 23
@@ H.1 Quality Assurance for Generations
- Lord Steven Regal won the World Television Championship.
- The Nasty Boys won the World Tag Team Championship.
@@
Your task is to generate a pair of sentences so that the provided claim can be entailed by the sentence pair. You must make sure that the claim can only be deduced by combining the information from the two sentences that contain unique information.
@@
We are creating a news article (one paragraph) in the style of The New York Times. We will give you a list of facts to use when writing your article. You must include all the facts in the list. Never state deduced facts or conclusions. The article should stick to the fact list pretty closely. Include as many sentences as needed to write each fact from the list of facts.